Richard has four times as many marbles as John. If Richard gave 18 to John they would have the same number. How many marbles has each?
Please tell me the exact steps to solve this, thanks :)

Answer by wuwei96815(245)

First, assign values to Richard and John's marbles based on the wording of the problem. We get:
x=John's marbles
4x=Richard's marbles

Next, write an equation based on the word clues in the probelm. We get:
4x-18=x

Finally solve the equation.
4x-18=x
4x-x=18
3x=18
x=6
4x=24

The main problem with word problems is converting the words to a math problem that can be solved. So, word problems are often primarily problems of interpretation.